In a family of 11 children, what is the probability that there will be more boys than girls?

1
Expert's answer
2021-11-30T15:53:10-0500

n=11

To get more boys than girls means at least 6 boys should be there for this condition

"P(Boy) = P(X) = \frac{1}{2} = 0.5 \\

P(Girl) = P(Y) = \frac{1}{2} = 0.5"

Probability for at least 6 boys:

"P(X≥6) = P(X=6) + P(X=7) + P(X+8) + P(X=9) + P(X=10) + P(X=11) \\

P(X=6) = C^{11}_6 (0.5)^6(1-0.5)^{11-6} = 0.2255859 \\

P(X=7) = C^{11}_7 (0.5)^7(1-0.5)^{11-7} = 0.1611328 \\

P(X=8) = C^{11}_8 (0.5)^8(1-0.5)^{11-8} = 0.0805664 \\

P(X=9) = C^{11}_9 (0.5)^9(1-0.5)^{11-9} = 0.0268555 \\

P(X=10) = C^{11}_{10} (0.5)^{10}(1-0.5)^{11-10} = 0.0053711 \\

P(X=11) = C^{11}_{11} (0.5)^{11}(1-0.5)^{11-11} = 0.0004883 \\

P(X≥6) = 0.50"

There is 50% probability that there is more boys than girls.